Consider the following information:

Net operating income under variable costing $25,000
Increase in inventory during the period 2,000 units
Fixed manufacturing overhead $50,000
Number of units produced during the period 10,000 units

Based on the above information, the net operating income under absorption costing is:

Group of answer choices

$15,000

$35,000

$75,000

$10,000
Answer #1

Fixed Manufacturing Cost Per Unit = Manufacturing Cost / Units Produced

= 50,000 /10,000 = $5 Per Unit

Increase in inventory During the Period = 2,000

Fixed OH in Ending inventory = 2,000*5 = 10,000

Net Operating income Under absorption Costing

= 25,000+10,000

= $35,000 (Answer)